Two jets leave an airport at the same time in opposite directions. The first jet is traveling at 377mph and the other at 275mph. How long will it take for the jets to be 2,065 miles apart?
Answer by mananth(16949)

I jet speed = 377 mph
II jet speed = 275 mph
Opposite directions
so add up speeds = 377+275= 652 mph
Distance = 2065 miles
Time = d/r
time = 2065/652
Time = 3.17 hours
= 3 hours 10 minutes